The probability of an event $A$ happening in one trial is $0.4$. The probability that the event $A$ happens at least once in three independent trials is

  • A
    $0.94$
  • B
    $0.784$
  • C
    $0.90$
  • D
    $0.22$

Explore More

Similar Questions

If $P(A_1 \cup A_2) = 1 - P(A_1^c) P(A_2^c)$,where $c$ stands for complement,then the events $A_1$ and $A_2$ are

$A$ box contains $15$ black and $20$ white balls. What is the probability that when $2$ balls are selected,at most one is white?

Difficult
View Solution

In a race,the odds in favour of horses $A, B, C,$ and $D$ are $1:3, 1:4, 1:5,$ and $1:6,$ respectively. Find the probability that one of them wins the race.

$A$ bag contains $5$ white,$7$ black,and $4$ red balls. Three balls are drawn from the bag at random. The probability that all the three balls are white is:

If $P(A) = \frac{1}{2}$,$P(B) = \frac{1}{3}$ and $P(A \cap B) = \frac{1}{4}$,then $P(B/A) = $

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo